OpenStudy (anonymous):

f(x)= 15-x^2 x greater than or equal to 0 f^-1 (x)= ?

OpenStudy (anonymous):

put \[x=15-y^2\] and solve for y

OpenStudy (anonymous):

you get \[y^2=15-x\] and so \[y=\pm\sqrt{15-x}\] but since the domain of your original function was positive numbers your answer is \[f^{-1}(x)=\sqrt{15-x}\]

OpenStudy (anonymous):

Got it! Thanks